One side of a triangle is twice the shortest side. the third side is 3 feet more that the shortest side. The perimeter is 19 feet. Find all three sides.
Answer by mukhopadhyay(490)

Let the shortest side be x feet long
The length of the second side = 2x
The length of the third side = x+3
Perimeter is the sum of all three sides
Thus, x+2x+(x+3) = 19
=>4x+3 = 19
=>4x = 16
=>x = 4
The sides of the triangle are 4 feet, 8 feet(=2*4), and 7 feet(=4+3) long.